OpenStudy (anonymous):

find the value of the expression: (sum difference formulas) cos(sin^-1(1/3)-tan^-1(1/2)

OpenStudy (anonymous):

subtraction angle formula for cosine here

OpenStudy (anonymous):

\[\cos(a-b)=\cos(a)\cos(b)-\sin(a)\sin(b)\] put \[a=\sin^{-1}(\frac{1}{3}),b=\tan^{-1}(\frac{1}{2})\]

OpenStudy (anonymous):

oops i made a typo sorry

OpenStudy (anonymous):

\[\cos(a-b)=\cos(a)\cos(b)\color{red}+\sin(a)\sin(b)\]

OpenStudy (anonymous):

get \[=\cos(\sin^{-1}(\frac{1}{2}))\cos(\tan^{-1}(\frac{1}{2}))\color{red}+\sin(\sin^{-1}(\frac{1}{2}))\sin(\tan^{-1}(\frac{1}{2}))\]

OpenStudy (anonymous):

of course you still have a pile of work to do